OpenStudy (anonymous):

Use the substitution method to solve the following system of equations. 4x – y = 3 7x – 9y = –2

OpenStudy (anonymous):

4x – y = 3 --> transform into y=4x-3 7x – 9y = –2 y=4x-3 7x – 9y = –2 -- sub in 7x-9(4x-3)=-2 7x-36x+27=-2 -29x=-29 x=1 sub back in y=4(1)-3 y=1 solution x=1 y=1

hero (hero):

y = 4x - 3 y = (2+ 7x)/9 Set y = y 9(4x-3) = 2 + 7x 36x - 27 = 2 + 7x 29 = 29x x = 1 y = 1